OpenStudy (anonymous):

1 3 2 - - 1- 4 8 (With working out please so i can understand it :D)

OpenStudy (sasogeek):

what's the meaning of the numbers? where did they come from? what topic? what's the question?

OpenStudy (anonymous):

Matrix? trying to see if there is a solution?

OpenStudy (sasogeek):

it does look like a 2 x 3 matrix but idk what the question is...

OpenStudy (anonymous):

Yeah and I dont understand what this means - 1-

OpenStudy (anonymous):

1 3 2 - 1- 4 8

OpenStudy (sasogeek):

what should we do with it?

OpenStudy (anonymous):

Make it into one number ?

OpenStudy (anonymous):

still doesn't make sense... why are the numbers stack up like that?

OpenStudy (anonymous):

Its 2 1/4 - 1 3/8

OpenStudy (anonymous):

9/4-11/8=18/8-11/8=(18-11)/8=7/8